Month of March, 2019:

                  Total deliveries made                                        278     

                  Shipments delivered late                                    16

                  Transit time for 12 sample shipments                 3, 3, 5, 4, 6, 5, 6, 4, 6, 7, 5, 4

                  Total units ordered                                            865,000

                  Total units shipped                                            830,400

                  Total units lost or damaged                                   4,800

                  March 31 inventory level                                    $849,000

                  March monthly sales                                         $2,575,000

Transportation Performance Analysis

What percentage of shipments were delivered on time in March 2019?

What was the average transit time across the sample of 12 deliveries?

Fulfillment Performance Analysis

What was the unit fill rate percentage for March 2019?

What was the unit damage rate percentage for March 2019?

Inventory Performance Analysis

What was the inventory turnover for March 2019?

How many days on hand of inventory were held in March 2019?
Answer #1

1. Total deliveries on time = 278- late deliveries = 278-16 = 262

%age of on time deliveries = 262/278 = 94.245

2. Average transit time = (3+3+5+4+6+5+6+4+6+7+5+4) / 12 = 4.83

3. Fill rate = Units Shipped / Units ordered = 830400 / 865000 = 96%

4. Unit damage rate = Damaged units / Units shipped = 4800 / 830400 =0.58%

5. Inventory turnover = March monthly sales / Inventory = $2,575,000 / $849,000 = 3.03

6. Days of inventory at hand = Number of days in March / Inventory turnover ratio = 31 / 3.03 = 10.23 days

Alternatively, daily sales =$2,575,000/31= 83064.52

Days of inventory at hand = Inventory / Daily sales = $849,000/ 83064.52 = 10.23 days